Svd on call shift summary l.jpg
This presentation is the property of its rightful owner.
Sponsored Links
1 / 30

SVD on-call shift summary PowerPoint PPT Presentation

SVD on-call shift summary 05/04/21 – 05/04/28 K.Hara,Y.Fujiyama,S.Ono There is no big problem. Troubles(1) error messages : http://belle.kek.jp/group/svd/ushiroda/svd/daq/elist.html 04/23 18:13:37 command STANDBY(28138):Rejected by SVD

Download Presentation

SVD on-call shift summary

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Svd on call shift summary l.jpg

SVD on-call shift summary

05/04/21 – 05/04/28

K.Hara,Y.Fujiyama,S.Ono

There is no big problem.

4/21-4/28, K.Hara,Y.Fujiyama


Troubles 1 l.jpg

Troubles(1)

error messages : http://belle.kek.jp/group/svd/ushiroda/svd/daq/elist.html

  • 04/23

    • 18:13:37 command STANDBY(28138):Rejected by SVD

      • -> checked the status window -> everything looked OK -> send command STANDBY again -> OK

      • I think that it was too early to send command.

  • 04/24

    • 17:58:28 ERROR from [SVD] PSON failed in sending “INIT(VA ON)” to TTM

      • -> check the status window -> TTM busy -> restart HD -> start new run

    • 18:10:54 FATAL from [SVD] PC DAQ fatal: PPCI card null read(0x22) from: SVD SP13 minor 0

      • -> start new run -> OK

4/21-4/28, K.Hara,Y.Fujiyama


Troubles 2 l.jpg

Troubles(2)

  • 04/28 5:00 Called by expert shift

    • FATAL from [SVD] [Relay process(fbdaq side)] Unexpected termination of HD in SVDlocal side.

      • -> Expert shift restart SVD x2. -> OK

      • known problem

    • Kawasaki-san who was BCG-shift in time told me this problem.

      • PC12 was hung up

4/21-4/28, K.Hara,Y.Fujiyama


Radiation dose l.jpg

Radiation Dose

you can get figure from http://belle.kek.jp/~tajimao/public/background/shift.gif

Total dose 379.350 kRad

Week dose 6.322 kRad

4/21-4/28, K.Hara,Y.Fujiyama


Leak current l.jpg

Leak current

It seemed that there is no a problem.

~svddaq/monitor/a.ps or http://belle.kek.jp/group/svd/monitor/hvmon.ps.gz

4/21-4/28, K.Hara,Y.Fujiyama


Local run l1 5 database l.jpg

Local Run / L1.5 database

On ill hybrid page

05/04/21

  • Pedestal of H139 seems to recover to reference.

4/21-4/28, K.Hara,Y.Fujiyama


Local run l1 5 database7 l.jpg

Local Run / L1.5 database

05/04/24

05/04/21

  • Total noise of H16-1,2 is higher than reference.

  • It has a peak on about 250ch after 24th.

4/21-4/28, K.Hara,Y.Fujiyama


Local run l1 5 database8 l.jpg

Local Run / L1.5 database

05/04/20

05/04/22

  • Total noise of H20(~250ch) have peak after 22th.

4/21-4/28, K.Hara,Y.Fujiyama


Local run l1 5 database9 l.jpg

Local Run / L1.5 database

On ill hybrid page

05/04/21

  • Total noise of H44(~510ch) seems to recover to reference.

4/21-4/28, K.Hara,Y.Fujiyama


Local run l1 5 database10 l.jpg

Local Run / L1.5 database

05/04/21

05/04/26

  • Gain of H158-3 was faint until 24th.

  • But, recover to reference after 26th.

4/21-4/28, K.Hara,Y.Fujiyama


Other comments l.jpg

Exp 43, Run 181

Exp 43, Run 176

Other Comments

  • For Beam vertex (r-phi) resol at SVD for Bhabhas, It seems that there is a peak at -0.009cm.

  • What is the cause of the problem?

4/21-4/28, K.Hara,Y.Fujiyama


Pedestal of hybrid all ref ps gz l.jpg

Pedestal of hybrid (all_ref.ps.gz)

4/21-4/28, K.Hara,Y.Fujiyama


Pedestal of hybrid l.jpg

Pedestal of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Total noise of hybrid l.jpg

Total noise of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Total noise of hybrid15 l.jpg

Total noise of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Total noise of hybrid16 l.jpg

Total noise of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Total noise of hybrid17 l.jpg

Total noise of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Total noise of hybrid18 l.jpg

Total noise of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Total noise of hybrid19 l.jpg

Total noise of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Gain of hybrid l.jpg

Gain of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Gain of hybrid21 l.jpg

Gain of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Gain of hybrid22 l.jpg

Gain of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Gain of hybrid23 l.jpg

Gain of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Gain of hybrid24 l.jpg

Gain of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Gain of hybrid25 l.jpg

Gain of hybrid

4/21-4/28, K.Hara,Y.Fujiyama


Gain of hybrid26 l.jpg

Gain of hybrid

  • I listed the local run results which are different from references.( faint, hot channel, etc)

4/21-4/28, K.Hara,Y.Fujiyama


Local run l1 5 database 1 l.jpg

Local Run / L1.5 database(1)

Local run results are available at

/fs00/svddaq/va_calib or http://belle.kek.jp/group/svd/ushiroda/svd/localrun

<0504211006>

Total noise : H16 140~470ch higher than ref.

Noise : H20-250ch,H205-150,200,460ch recover

H26-260ch,H48-370ch,H92-325ch up

Gain : H8-1 unsettled?,H139-50ch up,H27-0,H125-145ch recover

H19-250,370,460ch,H35-0,H51-25,125ch,H82-145ch,H135-50ch,

H107-100,390,480ch,H195-310ch down

-- Most of total noise were lower than ref. yesterday. But, they recovered today.

<0504221005>

Total noise : H16 140~470ch,H127-125ch same as yesterday,H20-260ch up

Noise : H16 160~380ch,H20-260ch,H104-350ch,H113-425ch,

H157 150~225ch,H199 up

H26-260ch,H48-370ch,H205-150,200,460ch same as yesterday

H92-325ch recover, H16 160~380ch a little higher than ref.

Gain : H8-1,H19-250,370,460ch,H168-1 recover,H195-160ch up,

H91-0,H107 100~130,480~512ch,H124-50ch,H148-125ch down,

H27-0,H125-145ch same as yesterday

4/21-4/28, K.Hara,Y.Fujiyama


Local run l1 5 database 2 l.jpg

Local Run / L1.5 database(2)

<0504231115>

Total noise : H16 140~470ch,H20-260ch same as yesterday

H70,71,78,79,82,86,144,145,146,147,156,157,202,203 higher than ref.

H127-125ch recover,H105-0ch up

Noise : H16 160~380ch,H20-260ch,H26-260ch,H104-350ch,H113-425ch,

H157 150~225ch same as yesterday

H48-370ch,H118-480ch,H195-160ch down

H92-325ch,H105-0ch,H145-270ch up

Gain : H19-2,3 down,H207-90ch up,H139-0,H164-0 unsettled

H8-1,H27-0,H35-0,H91-0,H195-160,310ch same as yesterday

H51-0,H82-145ch,H124-50ch,H148-125ch recover

<0504241006>

Total noise : H0,30,38,50,94,96,114,164,205 lower than ref.

H20-260ch,H146 same as yesterday,H16-250ch,H125-360ch up

H70,71,78,79,82,86,144,145,147,156,157,H105-0ch recover

4/21-4/28, K.Hara,Y.Fujiyama


Local run l1 5 database 3 l.jpg

Local Run / L1.5 database(3)

Noise : H26-260ch,H113-425ch,H199 same as yesterday

H125-360ch,H131-360ch,H173-140ch,H195-160ch up

H92-325ch,H104-350ch,H118-480ch down

H145-270ch,H157 150~225ch recover

Gain : H107,H139-0,H141 unsettled

H8-1,H19-2,3,H27-0,H35-0,H82-145ch,H91-0,H124-50ch,

H148-125ch,H164-0,H207-90ch same as yesterday

H195-310ch up,H51-100ch,H179-240ch,H195-160ch down

<0504260708>

Total noise : H0,H50,H97-240ch up,H113-40ch down

H16,20,30,70,71,78,79,82,86,94,105,114,144,145,147,156,157,169,

202,203 same as yesterday, H96,H146 recover

H177,H184 lower than ref.

Noise : H16,20,26,92,105,125,145,157,173,199 same as yesterday

H97-240ch,H104-350ch up,H113-420ch,H118-490ch,H134-190ch down

H131,195 recover

4/21-4/28, K.Hara,Y.Fujiyama


Local run l1 5 database 4 l.jpg

Local Run / L1.5 database(4)

Gain : H8-1 unsettled,H158-280ch,H195-160ch down,

H207 same as yesterday,H164,H179 recover,

H51-0,H59-0,3,H89-420ch,H107-1,2,H139-0 faint

<0504271009>

Pedestal : H53-1 down

Total noise : H16-260ch,H20-260ch,H32-1,H105-75ch,H173-125ch down

H97-240ch recover,H64-130ch,H71-250ch,H94-130ch,H101-0ch up,

-- Most of total noise of hybrid are lower than ref.

Noise : H16-260ch,H20-260ch,H173-125ch down,H71-250ch up,

H97-240ch recover

Gain : H19-2,3,H35-0,H51-0,195-160ch recover,H59-0,H158-280ch up,

H89-420ch,H164,H179-240ch same as yesterday,

H179-0 faint

4/21-4/28, K.Hara,Y.Fujiyama


  • Login